### v2

# Model X-ray : Detecting Backdoored Models via Decision Boundary

Abstract.
Backdoor attacks pose a significant security vulnerability for deep neural networks (DNNs), enabling them to operate normally on clean inputs but manipulate predictions when specific trigger patterns occur. In this paper, we consider a practical post-training scenario backdoor defense, where the defender aims to evaluate whether a trained model has been compromised by backdoor attacks. Currently, post-training backdoor detection approaches often operate under the assumption that the defender has knowledge of the attack information, logit output from the model, and knowledge of the model parameters, limiting their implementation in practical scenarios.
In contrast, our approach functions as a lightweight diagnostic scanning tool offering interpretability and visualization. By accessing the model to obtain hard labels, we construct decision boundaries within the convex combination of three samples. We present an intriguing observation of two phenomena in backdoored models: a noticeable shrinking of areas dominated by clean samples and a significant increase in the surrounding areas dominated by target labels. Leveraging this observation, we propose Model X-ray, a novel backdoor detection approach based on the analysis of illustrated two-dimensional (2D) decision boundaries. Our approach includes two strategies focused on the decision areas dominated by clean samples and the concentration of label distribution, and it can not only identify whether the target model is infected but also determine the target attacked label under the all-to-one attack strategy. Importantly, it accomplishes this solely by the predicted hard labels of clean inputs, regardless of any assumptions about attacks and prior knowledge of the training details of the model. Extensive experiments demonstrated that Model X-ray has outstanding effectiveness and efficiency across diverse backdoor attacks, datasets, and architectures. Besides, ablation studies on hyperparameters and more attack strategies and discussions are also provided.
Deep Learning, Backdoor Detection, Decision Boundary
$\dagger$ The corresponding author. E-mail: zhang_jie@cfar.a-star.edu.sg
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681075isbn: 979-8-4007-0686-8/24/10ccs: Security and privacyccs: Computing methodologies Machine learning
1. Introduction
Figure 1. Comparison of the decision boundaries between the clean model and the backdoored model (taking BadNets (Gu et al., 2017) as an example, and the target label is “airplane”) on the CIFAR-10 dataset.
Despite the remarkable success of DNNs, recent studies (Gu et al., 2017; Li et al., 2021a; Zeng et al., 2021; Wang et al., 2022; Liu et al., 2018; Doan et al., 2021; Bagdasaryan and Shmatikov, 2021) have unveiled a significant security vulnerability for DNNs against backdoor attacks, which can contaminate DNNs, enabling them to operate normally on clean inputs but manipulate predictions when specific patterns (i.e., “trigger”) occur. Backdoor attacks primarily fall into two categories: data-poisoning attacks (such as BadNets (Gu et al., 2017), SSBA (Li et al., 2021a), Low Frequency (Zeng et al., 2021), and BPP (Wang et al., 2022)) and model-modification attacks (such as TrojanNN (Liu et al., 2018), LIRA (Doan et al., 2021), and Blind (Bagdasaryan and Shmatikov, 2021)). These attacks pose a substantial threat to safety-critical and security-sensitive applications of DNNs, including but not limited to face recognition (Parkhi et al., 2015), biomedical diagnosis (Esteva et al., 2017), and autonomous driving (Redmon et al., 2016). To mitigate the threat of backdoor attacks, numerous defense methods are emerging to establish a comprehensive pipeline for backdoor defense. This pipeline can be applied at various stages, including the training, post-training, and deployment stages (refer to Fig. 2).
Backdoor defense during both the training and deployment stages (Chen et al., 2018; Tran et al., 2018; Gao et al., 2019; Zeng et al., 2021; Liu et al., 2023) typically necessitates access to training data or inference data. In this paper, we consider the more practical post-training scenario, where the defender aims to evaluate whether a trained model (e.g., Model Zoo that provides pre-trained models (zoo, 2018)) has been compromised by backdoor attacks, when and many post-training defenses assume the defender independently possesses a small set of clean, legitimate samples. However, current post-training detection methods hold too strong assumptions that the defender has knowledge of the attack information, the logit output from the model (Chen et al., 2019; Xu et al., 2021), and knowledge of the model parameters (Wang et al., 2019; Liu et al., 2019; Fu et al., 2023; Wang et al., 2023), limiting their implement in practical scenarios.
Fortunately, recent work by (Somepalli et al., 2022) has demonstrated that we can visualize the model’s decision boundary solely using prediction labels. Leveraging this technique, we have identified a discernible distinction between the decision boundaries of the clean model and the backdoored model. As illustrated in Fig. 1, we use BadNets (Gu et al., 2017) as an example of backdoor attacks. We observe that the decision boundaries of backdoor models exhibit a noticeable reduction in the regions dominated by three clean samples, and significant surrounding area are dominated by the attack target label. Importantly, this phenomenon is applicable across various backdoor attacks on different datasets (see Fig. 4). That is to say, we can leverage the phenomenon of anomalous decision boundaries to distinguish backdoored models. As claimed in previous work (Wang et al., 2019), backdoor attacks build a shortcut leading to the target label, which we explain cause the above encircling phenomena. Besides, trigger samples are more robust against distortions (Rajabi et al., 2023), causing the large regions than that of clean samples. In a nutshell, the visualized 2D decision boundary can be served as an illustration for these conjectures.
Based on the intriguing phenomenon, drawing an analogy to X-rays in disease diagnosis, we propose Model X-ray  as a novel backdoor detection approach through the analysis of illustrated 2D decision boundaries. Specifically, we designate two metrics to evaluate the degree of the closeness of the decision boundary: 1) Rényi Entropy (RE) (Rényi, 1961) calculated on the probability distribution of each prediction area and 2) Areas Dominated by triple samples (ATS), e.g., the total areas of “frog”, “ship”, and “dog” in the Fig. 1. Furthermore, if only one label is infected, we can determine the target label by the prediction of the largest area of the decision boundary, e.g., the target label is “airplane” in the right of Fig. 1. In other words, Model X-ray can not only identify backdoored models but also determine the target attacked label under all-to-one attacks. Importantly, Model X-ray accomplishes this only by the predicted hard labels of clean inputs from the model, regardless of any assumptions about attacks such as the trigger patterns and training details. The visualized 2D decision boundary offers a novel perspective to understand the behavior of the model, providing both visualization and interpretability. Through analysis of the decision boundary, Model X-ray can function as a lightweight diagnostic scanning tool, complementing other defense methods and aiding in defense pipelines. Extensive experiments demonstrate that Model X-ray performs better than current methods across various backdoor attacks, datasets, and model architectures. In addition, some ablation studies and discussions are also provided.
Our contributions can be summarized as follows:
  • We present a noteworthy observation: there exists a distinction between clean models and backdoored models as visualized through 2D decision boundaries (Somepalli et al., 2022). And we have visualized the decision boundaries of various backdoor attack models.
  • We propose Model X-ray  which detects the backdoored model solely by predicted hard labels of clean inputs from the model, regardless of any assumptions about backdoor attacks. Besides, Model X-ray can determine the target attacked label if the attack is all-to-one attack.
  • We evaluate Model X-ray using four combinations of datasets and architectures (including CNNs and ViTs) and seven backdoor attacks with diverse trigger types. Additionally, we assess Model X-ray on different backdoor attack label mapping strategies and varying poisoning rates. Extensive experiments demonstrate that Model X-ray has outstanding effectiveness and efficiency.
2. Related Work
2.1. Backdoor Attacks
The target of backdoor attacks is training an infected model $\hat{M}$ with parameters $\theta$ by:
(1) $\begin{array}[]{r}\theta=\arg\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\mathcal{L}(\hat{M}(x;\theta),y)\\ +\mathbb{E}_{\left(\hat{x},y_{t}\right)\sim\mathcal{\hat{D}}}\mathcal{L}(\hat{M}(\hat{x};\theta),y_{t}),\end{array}$
where $\mathcal{D}$ and $\mathcal{\hat{D}}$ denote the benign samples and trigger samples, respectively. $\mathcal{L}$ denotes the loss function, e.g., cross-entropy loss. The infected model functions normally on benign samples but yields a specific target prediction $y_{t}$ when presented with trigger samples $\hat{x}$. Backdoor attacks can be achieved by data poisoning and model modification, and we briefly introduce some related methods below.
Data poisoning-based backdoor attacks primarily revolve around crafting trigger samples. Notably, BadNets (Gu et al., 2017) was a pioneering work that highlighted vulnerabilities of DNNs by employing visible squares as triggers. Afterward, various other visible trigger techniques have been explored: Blended (Chen et al., 2017) employs image blending to create trigger patterns, SIG (Barni et al., 2019) utilizes sinusoidal strips as triggers, and Low Frequency (LF) (Zeng et al., 2021) explores triggers in the frequency domain. Simultaneously, other research endeavors focus on achieving imperceptibility of the trigger patterns, including BPP (Wang et al., 2022) based on image quantization and dithering, WaNet (Nguyen and Tran, 2021) founded on image warping, and SSBA (Li et al., 2021a) achieved by image steganography. During the training stage, the attacker can leverage different poisoning ratios to balance the attack ability and performance degradation.
Apart from data poisoning-based attacks, there are some backdoor attacks that employ model modification techniques. TrojanNN (Liu et al., 2018) first proposes to optimize the trigger to ensure that the crucial neurons can attain their maximum values, LIRA (Doan et al., 2021) formulates malicious function as a non-convex, constrained optimization problem to learn invisible triggers through a two-stage stochastic optimization procedure, and Blind (Bagdasaryan and Shmatikov, 2021) modifies the training loss function to enable the model to learn the malicious function.
2.2. Backdoor Defenses
Figure 2. The pipeline of the backdoor defense.
As Fig. 2 illustrates, pipelines for backdoor defense mechanisms can be categorized into three phases: during training, post-training, and after deployment. Each phase implies distinct defender roles and capabilities.
Backdoor defenses during model training aim to detect and remove poisoned data from the training set (Chen et al., 2018; Tran et al., 2018; Tang et al., 2021) or to enhance training robustness against data poisoning (Li et al., 2021b). Backdoor defenses after deployment aim to detect trigger inputs during inference and attempt to mitigate the malicious prediction. For example, STRIP(Gao et al., 2019) perturbs an input sample by overlapping with numerous benign samples and uses the ensemble predictions for detection. FreqDetector (Zeng et al., 2021) leverages artifacts in the frequency domain to distinguish trigger samples from clean samples. Besides, some methods (Liu et al., 2023; Rajabi et al., 2023; Guo et al., 2023) conduct detection based on robustness against data transformations between benign and trigger samples.
Comparably, post-training backdoor detection is model-level detection. Neural Cleanse (Wang et al., 2019) is the first post-training detection through anomaly analysis on the reversed trigger patterns. However, it requires access to the model’s inner information like parameters and gradients, which is also the limitation of other subsequent methods (Wang et al., 2019; Liu et al., 2019; Wang et al., 2020; Fu et al., 2023; Wang et al., 2023). Differently, detection work in black-box scenarios is extremely challenging (Chen et al., 2019; Xu et al., 2021; Guo et al., 2021; Dong et al., 2021), e.g. MNTD trains a meta-classifier based on features extracted from a large set of shadow models. However, its success heavily relies on the generalization capability of the attack settings from the shadow models to the actual backdoored models. Besides, it requires the soft label generated by the target model. MM-BD (Wang et al., 2023) leverags maximum margin statistics of each class and unsupervised anomaly detection on classifier output landscapes.
2.3. Decision Boundary of Deep Neural Networks
Most previous works depict decision boundaries by adversarial samples (He et al., 2018; Khoury and Hadfield-Menell, 2018) or sensitive samples (He et al., 2019). These methods are pivotal in identifying and understanding the contours of decision boundaries, as adversarial and sensitive samples are typically positioned along these critical junctures in the model’s decision-making process. However, obtaining these special samples requires access to the target model. Fortunately, Zhang et al. (Zhang et al., 2017) find that decision boundaries not only manifest near the data manifold but also within the convex hull created by pairs of data points.
Leveraging this understanding, Somepalli et al. (Somepalli et al., 2022) introduce an innovative approach that utilizes only clean samples to map out the decision boundary to investigate reproducibility and double descent. Their method, which results in a 2D map, offers an intuitive and accessible means of visualizing decision boundaries. In this paper, we utilize this technique to detect backdoored models.
3. Preliminaries
3.1. Recap of the Decision Boundary in (Somepalli et al., 2022)
Here, we recap the methods for visualizing decision boundaries discussed in (Somepalli et al., 2022). As shown in Fig. 3 (left), we randomly choose three clean samples (also called triple samples) from the dataset $\mathcal{D}.$ For example, we select three images ($x_{1}$, $x_{2}$, $x_{3}$) of “frog”, “ship”, and “dog” from the CIFAR-10 dataset. Then, we can calculate two vectors $\overrightarrow{v_{1}}$ = $x_{2}-x_{1}$ and $\overrightarrow{v_{2}}$ = $x_{3}-x_{1}$, based on which we obtain the spanned space $\mathcal{V}$, i.e., $\mathcal{V}=span\{\overrightarrow{v_{1}},\overrightarrow{v_{2}}\}$, whose orthogonal basis and orthonormal basis are denoted as $\{\overrightarrow{\beta_{1}},\overrightarrow{\beta_{2}}\}$ and $\{\overrightarrow{e_{1}},\overrightarrow{e_{2}}\}$, respectively, where $\overrightarrow{\beta_{1}}=\overrightarrow{v_{1}}$ and $\overrightarrow{e_{1}}=\frac{\overrightarrow{\beta_{1}}}{\|\overrightarrow{\beta_{1}}\|}$ Next, we can obtain the projection of vector $\overrightarrow{v_{2}}$ in the direction of vector $\overrightarrow{e_{1}}$, i.e., $\operatorname{proj}_{\overrightarrow{e_{1}}}\overrightarrow{v_{2}}$ = ${\Braket{\overrightarrow{v_{2}},\overrightarrow{e_{1}}}}\cdot\overrightarrow{e_{1}}$ and get $\overrightarrow{\beta_{2}}$ by orthogonalizing $\overrightarrow{v_{2}}$ via Schmidt orthogonalization, i.e., $\overrightarrow{\beta_{2}}=$$\overrightarrow{v_{2}}-\operatorname{proj}_{\overrightarrow{e_{1}}}\overrightarrow{v_{2}}$. Similarly, we can acquire the projection of vector $\overrightarrow{v_{2}}$ in the direction of vector $\overrightarrow{e_{2}}$, i.e., $\operatorname{proj}_{\overrightarrow{e_{2}}}\overrightarrow{v_{2}}$ = ${\Braket{\overrightarrow{v_{2}},\overrightarrow{e_{2}}}}\cdot\overrightarrow{e_{2}}$. Finally, we obtain an orthonormal basis for the space, denoted as $\overrightarrow{e_{1}}$ and $\overrightarrow{e_{2}}$, along with the coordinates of points $x_{1}$, $x_{2}$, and $x_{3}$ within the plane. Namely, we acquire coordinates corresponding to the origin (0, 0) and the points specified by vectors $\overrightarrow{v_{1}}$ and $\overrightarrow{v_{2}}$, originating from the origin, i.e., (0, 0), ($\|\overrightarrow{v_{1}}\|$, 0), ($\operatorname{proj}_{\overrightarrow{e_{1}}}\overrightarrow{v_{2}}$, $\operatorname{proj}_{\overrightarrow{e_{2}}}\overrightarrow{v_{2}}$).
After representing the space, we can calculate the bounds on the X-axis and the Y-axis, extended by a factor of $\eta$ in both the positive and negative directions along the corresponding axes, serving as a means to control the expansion range of the coordinate system. In the previous work (Somepalli et al., 2022), $\eta$ is set as 1 to investigate reproducibility, while we set $\eta$ as 5 to obtain a wider range of the decision boundary (see Fig. 3). Moreover, we can also determine density $S$ by constructing the set of points with a quantity of $S^{2}$ within the bounded range of the coordinate system using a grid generation method. Larger $S$ means higher resolution. With $S^{2}$ points, we can conduct the reverse process to get their tensor presentation, which can be fed to the model to fetch the corresponding prediction. We adopt different colors for different predictions to get the final 2D decision boundary.
In the subsequent parts, all decision boundaries are visualized by the modified version (i.e., $\eta=5$ in the right of Fig. 3).
Figure 3. Visual examples of the decision boundary used in (Somepalli et al., 2022) (left) and in this paper (right).
3.2. Threat Model
In practice, acccess to training or inference sets is unavailable due to data privacy, ownership, and availability constraints. Therefore, in this paper, we only consider the post-training scenario detection.
While many post-training detection defenses typically have access to either the model’s weights(Wang et al., 2019; Liu et al., 2019; Wang et al., 2020; Fu et al., 2023; Wang et al., 2023) or the model’s logit output(Xu et al., 2021) for evaluation, our approach goes a step further by restricting access to the model. We only assume that the defender has the capability to independently gather a small set of clean data samples that cover all classes within the domain, a prerequisite upon which most post-training detectors depend. Moreover, we only need the hard label predictions of the target model.
4. Method
In this section, we first provide an intriguing observation on the decision boundary of clean models and backdoor models. Based on this, we designate two strategies for backdoor detection via the decision boundary. Finally, we showcase that we can determine the target attacked label, if only one single label is infected.
Figure 4. Visual examples of decision boundaries of the clean model and different backdoored models on CIFAR-10 and ImageNet-10.
4.1. An Intriguing Observation
As shown in Fig. 4, we provide the decision boundary of the clean model and different backdoored models (infected by BadNets (Gu et al., 2017), SSBA (Li et al., 2021a), LF (Zeng et al., 2021), BPP (Wang et al., 2022), TrojanNN (Liu et al., 2018), LIRA (Doan et al., 2021), and Blind (Bagdasaryan and Shmatikov, 2021)) on CIFAR-10 and ImageNet-10 dataset. We observe two key phenomena in backdoored models: (1) the noticeable shrinking of areas dominated by clean samples and (2) the significant increase in surrounding areas dominated by target labels. More visualized decision boundaries can be found in the supplementary material.
We explain this phenomenon may be the shortcut effect caused by backdoor attacks. In essence, clean models can still preserve the robustness of predicted labels when applying a linear transformation to samples in a considerably large magnitude. On the contrary, the presence of shortcuts to the target label in backdoor models results in changes in the predicted label when applying a minor linear transformation to samples, typically leading to the target attacked label. The shortcuts leading to the target attacked label in the backdoor model has been confirmed in previous research, that is, through optimization methods, smaller perturbations can be found to cause other labels to be misclassified as target labels (Wang et al., 2019). Afterward, Rajabi et al. (Rajabi et al., 2023) quantifies this effect by introducing the concept of a certified radius (Cohen et al., 2019), which estimates the distance to a decision boundary by perturbing samples with Gaussian noise with a predetermined mean and variance. Notably, trigger samples are observed to be relatively farther from a decision boundary compared to clean samples, which can support why the large region is dominated by injected prediction.
Figure 5. Illustration on calculation of RE and ATS.
4.2. Two Strategies for Backdoor Detection via the Decision Boundary
Based on two key phenomena, we have developed two corresponding strategies that focus on the decision areas dominated by clean samples and the concentration of label distribution, namely, based on Rényi Entropy (RE) and Areas dominated by Triple Samples (ATS).
4.2.1. Backdoor Detection based on Rényi Entropy
With the technique mentioned above, we can plot $N$ decision boundaries $\mathbf{B}=\{\mathcal{B}_{1},...,\mathcal{B}_{k},...,\mathcal{B}_{N}\}$, where $\mathcal{B}_{k}$ is plotted along the plane spanned by triple samples $T_{k}=(x_{1},x_{2},x_{3})_{k}$. Specifically, let $S_{k}=\{x_{ij}|(i,j)\in\mathcal{B}_{k}\}$ be the set of points in the $\mathcal{B}_{i}$, where $(i,j)$ is the coordinations of $x$ in $\mathcal{B}_{k}$. Then, we feed $S_{k}=\{x_{ij}|(i,j)\in\mathcal{B}_{k}\}$ to the target model $M$ to obtain the corresponding hard labels $L_{k}=\{l_{ij}|(i,j)\in\mathcal{B}_{k}\}$, which are further used to obtain the final colorful decision boundary $\mathcal{B}_{k}$ for evaluation.
Within a specific decision boundary $\mathcal{B}_{k}$, we calculate label probability distribution $\mathcal{P}_{k}=\{p_{1},...,p_{m},...,p_{n}\}$ for n-category classification:
(2) $\small p_{m}=\frac{A(l_{m})}{A(\mathcal{B}_{k})},$
[…]
(3) $\small RE(\mathcal{P}_{k})=H_{\alpha}(\mathcal{P}_{k})=\frac{1}{1-\alpha}\log\left(\sum_{m=1}^{n}p_{m}^{\alpha}\right),$
where $\alpha\geqslant 1$, and we set it as 10 by default. Based on RE, we propose a detection strategy called Ours-RE. Briefly, a large variance of $\{p_{1},...,p_{m},...,p_{n}\}$ will lead a low RE, meaning more gathered. As shown in Fig. 4, we find backdoored models hold much lower RE, which can be distinguished from the clean model in most cases.
4.2.2. Backdoor Detection based on Areas dominated by Triple Samples
In addition to RE, we define Areas dominated by Triple Samples (ATS) as the ratio of decision regions controlled by benign triple samples $T_{k}$ to entire decision regions:
(4) $\small ATS(\mathcal{B}_{k})={\frac{A(T_{k})}{A(\mathcal{B}_{k})}}={\frac{\sum_{x\in{(x_{1},x_{2},x_{3})}}{A(x)}}{A(\mathcal{B}_{k})}},$
[…]
Figure 6. The label probability distribution within decision boundaries of clean and backdoor models on CIFAR-10 and ImageNet-10, both of whose infected labels are 0.
4.3. Determine the Target Label
After detecting, if the attack is conducted by all-to-one strategy, defenders can further determine the target attacked label by identifying the label with an abnormally high probability in label probability distribution $\mathcal{P}_{k}=\{p_{1},...,p_{m},...,p_{n}\}$. For example, we plot decision boundaries of clean models and backdoor models infected by different backdoor attacks on CIFAR-10 and ImageNet-10 datasets. For each model, we plot 20 decision boundaries and calculate the average label probability. As shown in Fig. 6, the attacked target label (label “0” of both CIFAR-10 and ImageNet-10) exhibits an exceptionally high probability, even reaching 80% to 90% of the entire label probability distribution.
5. Experiment
5.1. Experimental Settings
Datasets and Architectures. The datasets include CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), GTSRB (Houben et al., 2013), and ImageNet-10 (imt, 2020), a subset of ten classes from ImageNet (Deng et al., 2009). Besides, we employ four different architectures: PreActResNet-18 (He et al., 2016), MobileNet-V3-Large (Howard et al., 2019), PreActResNet-34 (He et al., 2016), and ViT-B-16 (Dosovitskiy et al., 2020). These architectures encompass both Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) and span across various network sizes, including small, medium, and large networks.
Implementation Details. For the model to be evaluated, we plot decision boundaries by random samples triplet with expansion factor $\eta=5$ and density $S=100$, number of plots $N=20$. For the attack baselines, we evaluate our method against seven backdoor attacks, including BadNets (Gu et al., 2017), SSBA (Li et al., 2021a), LF (Zeng et al., 2021), BPP (Wang et al., 2022), TrojanNN (Liu et al., 2018), LIRA (Doan et al., 2021), and Blind (Bagdasaryan and Shmatikov, 2021). We follow an open-sourced backdoor benchmark BackdoorBench (Wu et al., 2022) for the training settings of these attacks and conduct all-to-one attacks by default. As shown in Table 1, the attacks in our experiments include both data poisoning-based attacks and model modification-based attacks, which contain diverse and complex trigger pattern types. In this paper, our focus is on post-training backdoor detection. We compare our approach with three post-training detection methods: Neural Cleanse (Wang et al., 2019), MNTD (Xu et al., 2021), and MM-BD (Wang et al., 2023). We utilize their official implementations (mnt, 2020; mmb, 2023) or implementations available in open-source benchmarks(Bac, 2023).
Table 1. The backdoor attacks involved in our evaluations have covered diverse trigger patterns.
[…]
Invisible
[…]
Evaluation Metrics. For clean models and models infected by 7 backdoor attacks, we trained 20 models using different initialization and random seeds. For the backdoored models, we select different attack target labels and conduct the single-label attack by default. Considering the computational cost, we adopted different data sets and corresponding common model architectures. Thus, we have $20+20\times 7=160$ models for each combination of dataset and architecture. In subsequent experiments, for each model to be evaluated, we calculate its average RE (see Eq. (3)) and ATS (see Eq. (4)) over $N=20$ decision boundary plots as indicators. We assume that defense mechanisms return a positive label if they identify a model as a backdoored model and then compute the Area Under Receiver Operating Curve (AUROC) to measure the trade-off between the false positive rate (FPR) for clean models and true positive rate (TPR) for backdoor models for a detection method.
Table 2. The performance of Model X-ray across different attacks, datasets, and architectures. The last two columns show the worst and the average performance among different attacks. The best results are in bold.
[…]
MNTD 0.525 0.665 0.568 0.565 0.568 0.623 0.705 0.525 0.603
MM-BD 1.000 0.847 0.882 0.805 0.860 0.953 0.697 0.697 0.863
PreActResNet-18 Ours-RE 0.995 1.000 0.812 0.762 0.740 1.000 0.919 0.740 0.890
[…]
MNTD 0.603 0.495 0.578 0.617 0.535 0.715 0.460 0.460 0.572
MM-BD 1.000 0.477 0.494 0.445 0.792 0.994 0.997 0.445 0.743
MobileNet-V3 Ours-RE 0.997 0.981 0.942 1.000 0.976 1.000 1.000 0.942 0.985
[…]
MNTD 0.625 0.490 0.540 0.528 0.540 0.813 0.538 0.490 0.582
MM-BD 0.626 0.552 0.977 0.557 0.618 0.957 0.633 0.552 0.703
PreActResNet-34 Ours-RE 1.000 1.000 1.000 0.832 0.746 0.979 0.819 0.746 0.911
[…]
MNTD 0.588 0.428 0.620 0.323 0.620 0.632 0.478 0.323 0.527
MM-BD 0.107 0.205 0.135 0.120 0.215 0.518 0.149 0.107 0.207
ViT-B-16 Ours-RE 0.956 0.860 0.835 0.913 0.725 1.000 0.863 0.725 0.879
[…]
(a) CIFAR-10
(b) GTSRB
(c) CIFAR-100
(d) ImageNet-10
Figure 7. The average RE ($\alpha=10$) for clean and backdoor models injected by seven backdoor attacks in CIFAR-10, CIFAR-100, GTSRB, and ImageNet-10 datasets. We observe that backdoor models have significantly smaller RE than clean models.
(a) CIFAR-10
(b) GTSRB
(c) CIFAR-100
(d) ImageNet-10
Figure 8. The average ATS ($t=0.5$) for clean and backdoor models injected by seven backdoor attacks in CIFAR-10, CIFAR-100, GTSRB, and ImageNet-10 datasets. We observe that backdoor models have significantly smaller ATS than clean models.
5.2. The Effectiveness of Model X-ray
As shown in Table 2, in most cases, Model X-ray outperforms the baseline methods across different backdoor attacks, datasets, and architectures. MNTD is difficult to generalize attack settings from the shadow models to the actual backdoored models. Neural Cleanse performs well in the majority of scenarios. However, occasional failures may arise when it incorrectly identifies a trigger for a clean model, leading to convergence in local optima. MM-BD demonstrates promising performance on small-scale architectures, but its performance drops significantly on larger architectures. In Fig. 7 and Fig. 8, we present visual illustrations of the average RE and ATS values for both clean and backdoored models. In most cases, a clear distinction is evident between clean and backdoored models. The ROC curves of Ours-RE and Ours-ATS can be found in the supplementary material.
Table 3. The performance under different attack strategies.
[…]
MNTD 0.525 0.419 0.503 0.487 0.535 0.518 0.466
MM-BD 1.000 0.571 0.006 0.081 0.007 0.448 0.671
Ours-RE 1.000 0.995 0.824 0.829 0.839 0.638 0.423
[…]
(a) PreActResNet18
(b) MobileNet-L
(c) PreActResNet34
(d) VGG19
Figure 9. The average RE ($\alpha=10$) for clean and backdoor models injected by seven backdoor attacks in CIFAR-10 on different architectures.
(a) PreActResNet18
(b) MobileNet-L
(c) PreActResNet34
(d) VGG19
Figure 10. The average ATS ($t=0.5$) for clean and backdoor models injected by seven backdoor attacks in CIFAR-10 on different architectures.
Figure 11. Decision boundaries under different attack strategies.
Besides the default all-to-one attack strategy, we consider attack strategies (Xiang et al., 2023) with arbitrary numbers of source classes each assigned with an arbitrary attack target class, including X-to-X attack, X-to-one attack, and one-to-one attack. We adopt different attack strategies to conduct BadNets on CIFAR-10. For each strategy, we train 10 models for evaluation. Table 3 shows that Model X-ray remains effective under different attack strategies, especially based on ATS (i.e., Ours-ATS). Although multi-target attacks lower the performance of the proposed method, we outperform the baseline methods by a large margin in most cases. Furthermore, we provide some visual examples of the corresponding decision boundary in Fig. 11. In X-to-one and one-to-one attacks, where the attack target is a single class, both Ours-RE and Ours-ATS achieve precise detection and identification of the target class. In X-to-X attack, where there are multiple classes for both source and attack targets, the performance of Ours-RE declines with an increasing number of attack target classes, which is acceptable. The computation of Ours-RE relies on the entropy of class labels, where it can still detect the presence of multiple attack target classes in the decision boundary, despite the performance drop. Furthermore, areas dominated by triple clean samples shrink, which explains why Ours-ATS achieves good performance in such scenarios.
5.3. Evaluations on Open-source Benchmarks
To mitigate the impact of incidental factors in our training, we also evaluated our method on the backdoored models pre-trained on an open-source benchmark (Wu et al., 2022). Specifically, we perform detection on pre-trained backdoored models injected with seven backdoor attacks across CIFAR-10, GTSRB, and CIFAR-100 datasets using the PreActResNet-18 architecture, which can be downloaded from Open-source benchmarks (Bac, 2023).
Given a target model $C_{\theta}$, Model X-ray map the model $C_{\theta}$ to a linearly separable space, defenders can make judgments through average RE and ATS based on a threshold $\gamma$ :
[…]
As shown in Fig. 9 and Fig. 10, for the same dataset (taking CIFAR-10 as an example), we find that the relationship of RE and ATS between clean and backdoor models exhibits consistency. This allows us to determine an estimated threshold $\bar{\gamma}$ based on a small set of models:
(6) $\bar{\gamma}=\frac{1}{N}\sum_{m=1}^{N}\arg\max_{\gamma\in\Gamma}\frac{2\times\left(\text{ precision }_{\gamma}\times\text{ recall }_{\gamma}\right)}{\left(\text{ precision }_{\gamma}+\operatorname{recall}_{\gamma}\right)}.$
Based on thresholds $\bar{\gamma}$ (e.g., for Ours-RE CIFAR-10: $0.873$, GTSRB: $2.040$, CIFAR-100: $1.194$; for Ours-ATS, CIFAR-10: $0.184$, GTSRB: $0.134$, CIFAR-100: $0.040$), the detection accuracy on CIFAR-10 is 87.5%, on GTSRB is 93.75% and on CIFAR-100 is 100%. The visualized decision boundaries can be found in the supplementary material.
5.4. The Efficiency of Model X-ray
Neural Cleanse and MM-BD necessitate access to the model’s parameters, and MNTD relies on logit outputs from the target model. Model X-ray detects the backdoored model solely by predicted hard labels of clean inputs from the model. In Table 4, we show the number of benign samples that the defender needs. Both Neural Cleanse and MNTD necessitate a certain proportion of benign data (e.g., 5% of the benign dataset) to complement their defense mechanisms, MM-BD does not require any clean data. Our method necessitates only three benign samples to plot a decision boundary, and with $N$ set to 20, only 60 clean samples are required, which is already sufficient to ensure the effectiveness of our detection.
In addition, we compare the average inference time of each method in Table 5. The experiment is conducted on one NVIDIA RTX A6000. Specifically, Neural Cleanse requires a trigger reverse engineering optimization process for each class, MM-BD also requires a margin statistical process to obtain a maximum margin statistic for each class, and MNTD requires preparation that generates a large set of shadow models (1024 clean models and 1024 attack models) to train a meta-classifier. In contrast, our method eliminates the need for any optimization or training processes, making it a versatile plug-and-play solution that functions as a lightweight diagnostic scanning tool.
Table 4. Benign samples required for different methods.
Method CIFAR-10 GTSRB CIFAR-100 ImageNet-10
Neural Cleanse 2500 1332 2500 473
MNTD 2500 1332 2500 473
MM-BD 0 0 0 0
Ours 60 60 60 60
Table 5. The average inference time(sec) for different methods. $\dagger$ means the training time(sec).
Method Neural Cleanse MNTD $\dagger$ MNTD MM-BD Ours
CIFAR-10 243.4 44268.6 0.06 75.2 36.5
GTSRB 628.5 53409.0 0.05 334.8 34.6
CIFAR-100 2431.7 46680.9 0.06 829.5 36.0
ImageNet-10 1471.0 73632.2 1.5 414.4 112.3
Figure 12. The influence of the number of plots $N$ and point density $S$.
Figure 13. The influence of the parameters $\alpha$ and $t$.
Figure 14. The influence of the poisoning ratio.
5.5. Ablation Study
The Influence of the Hyper-parameters. $N$ is the number of decision boundary plots and $S$ is the density of decision boundaries, which are critical to the evaluation efficiency. Here, we investigate Model X-ray’s performance under fixed $N=20$ with $S$ ranging from 60 to 140 and under fixed $S=100$ with $N$ ranging from 5 to 40. Fig. 12 shows that lower $N$ and $S$ will slightly degrade the performance of Model X-ray on CIFAR-10, which is still acceptable. Besides, we investigate the impact of parameters in two indicators, i.e., $\alpha$ in RE and $t$ in ATS. As shown in Fig. 13, different $\alpha$ has a neglectable effect on Ours-RE, while $t$ larger than 0.5 is better for Ours-ATS.
The Influence of the Poisoning Ratio. In the above experiment, we set the poisoning ratio as 10% by default. Here, we further evaluate our method against data-poisoning attacks under different poisoning ratios (1%, 5%, 10%, and 20%) on CIFAR-10 dataset. As shown in Fig. 14, as the poisoning ratio increases, our approach becomes more effective, indicating that the phenomenon of anomalous decision boundaries in the backdoor models becomes more pronounced. For low ratios like 1%, the attack ability for some attacks degrades, wherein the poorer performance is understood.
Figure 15. Decision boundaries of Blended (Chen et al., 2017) and WaNet (Nguyen and Tran, 2021).
6. Discussion
Special Cases. We find that Ours-AST can distinguish between the backdoored WaNet model and a clean model. Interestingly, the AST of WaNet is larger than that of clean models (see Fig. 15), possibly because WaNet enhances the robustness of clean samples. However, Blended (Chen et al., 2017) can bypass Ours-RE, likely because blending the trigger pattern with clean samples does not create shortcuts due to model redundancy.
7. Conclusion
In this paper, we observe a distinction between clean and backdoored models through visualized 2D decision boundaries. Based on this, we propose Model X-ray, a novel black-box hard-label post-training backdoor detection approach.
Acknowledgements
This work is supported in part by the Natural Science Foundation of China under Grant 62121002, U20B2047, U2336206, 62102386 and 62372423 and by Anhui Province Key Laboratory of Digital Security, and by National Research Foundation, Singapore and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN), and Infocomm Media Development Authority under its Trust Tech Funding Initiative (No. DTC-RGC-04), and Singapore Ministry of Education (MOE) AcRF Tier 2 under Grant MOE-T2EP20121-0006.
Supplementary Materials for Model X-ray : Detecting Backdoored Models via Decision Boundary
Appendix A Implementation Details
A.1. Details of Datasets
Table 6 shows the details of four datasets for evaluation.
Table 6. The datasets for evaluation
Dataset Classes Image Size Training Data Test Data
CIFAR-10 10 3×32×32 50,000 10,000
GTSRB 43 3×32×32 39,209 12,630
CIFAR-100 100 3×32×32 50,000 10,000
ImageNet-10 10 3×160×160 9,469 3,925
A.2. Details of Training Configurations
Following an open-sourced backdoor benchmark (Wu et al., 2022), we use the same training configuration to train both clean and backdoored models. These details are shown in Table 7.
Table 7. The training configurations for both clean and backdoored models.
Configurations CIFAR-10 GTSRB CIFAR-100 ImageNet-10
Model PreActResNet-18 MobileNet-V3-Large PreActResNet-34 ViT-B-16
Optimizer SGD SGD SGD SGD
SGD Momentum 0.9 0.9 0.9 0.9
Batch Size 128 128 128 128
Learning Rate 0.01 0.01 0.01 0.01
Scheduler CosineAnnealing CosineAnnealing CosineAnnealing CosineAnnealing
Weight Decay 0.0005 0.0005 0.0005 0.0005
Epochs 50 50 50 50
A.3. Details of Attacks Configurations
For backdoored models, we conduct all-to-one attacks by randomly choosing the attack target label. For poisoning-based attacks, we set the default poisoning ratio as 10%, for model modification-based attacks, we adopt its default implementation in (Wu et al., 2022).
The attacks involved in our evaluations contain diverse complex trigger pattern types, including patch, invisible, and input-aware triggers. Fig. 16 shows some visual examples of trigger samples generated by diverse backdoor attacks involved in our evaluations. Table 8 shows attack ability for diverse backdoor attacks on different datasets and architectures, indicating that all the attacks are effective.
Table 8. The average accuracy (ACC) and attack success rate (ASR) for different backdoored models on different datasets and architectures.
Dataset Architecture Clean BadNets SSBA LF BPP TrojanNN LIRA Blind
ACC ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
CIFAR-10 PreActResNet-18 0.925 0.913 0.963 0.902 0.955 0.906 0.969 0.910 0.964 0.917 0.999 0.648 0.806 0.845 0.950
GTSRB MobileNet-V3 0.904 0.891 0.922 0.905 0.853 0.905 0.995 0.889 0.823 0.908 0.999 0.129 0.951 0.787 0.791
-Large
CIFAR-100 PreActResNet-34 0.712 0.664 0.915 0.692 0.968 0.682 0.954 0.660 0.985 0.695 0.999 0.158 0.984 0.562 0.999
ImageNet-10 ViT-B-16 0.985 0.983 0.997 0.984 0.993 0.869 0.804 0.963 0.958 0.984 0.998 0.942 0.965 0.982 0.999
Figure 16. Some visual examples of trigger samples for different backdoor attacks.
Figure 17. The ROC curves of Ours-RE [top] and Ours-ATS [bottom] against seven backdoor attacks on CIFAR-10, GTSRB, CIFAR-100, and ImageNet-10.
A.4. Details of the Defense Baselines
As shown in Fig. 17, we present the ROC curves of our methods. For Neural Cleanse (Wang et al., 2019), we adopt its default implementation in (Wu et al., 2022). For MNTD (Xu et al., 2021), we adopt its official implementation (mnt, 2020) to train a meta-classifier based on features extracted from a large set of shadow models (1024 clean models and 1024 attack models) for each combination of dataset and architecture, it means ${2048\times 4=8192}$ shadow models in total. For MM-BD, we adopt its official implementation in (mmb, 2023).
Appendix B The Influence of Expansion Factor $\eta$ for Plotting Decision Boundary
In the above experiment, we set the value for the expansion factor $\eta$ to 5 by default, which is a balanced choice. A larger $\eta$ may imply greater computational overhead, while a smaller $\eta$ may not adequately exhibit the effects of backdoor attacks.
Here, we further evaluate Ours-RE by setting the expansion factor $\eta$ to 3, 5 and 8 on the CIFAR-10 dataset. As shown in Table 1, we observed anomalous decision boundary phenomena in certain attacks occurring over a larger range, such as in LF, BPP, and TrojanNN. Therefore, choosing a larger value for $\eta$ can enhance the performance of Ours-RE.
Table 9. The influence of expansion factor $\eta$.
[…]
N=5 0.995 1.000 0.812 0.762 0.740 1.000 0.919 0.890
[…]
Appendix C More evaluation results
Evaluations on Open-source Benchmarks
Based on thresholds $\bar{\gamma}$ (e.g., for Ours-RE CIFAR-10: $0.873$, GTSRB: $2.040$, CIFAR-100: $1.194$; for Ours-ATS, CIFAR-10: $0.184$, GTSRB: $0.134$, CIFAR-100: $0.040$), the detection accuracy on CIFAR-10 is 87.5%, on GTSRB is 93.75% and on CIFAR-100 is 100%. As shown in Fig. 18, Model X-ray consistently identifies anomalies in the decision boundaries that three samples are encircled by a large area of the target label, demonstrating precise detection of backdoored models and determine the attack target labels.
Figure 18. Evaluation results for backdoored models on CIFAR-10, GTSRB, and CIFAR-100 using PreActResNet-18 pre-trained on an open-source benchmark (Bac, 2023).
Appendix D More Visual Examples of Decision Boundaries for Backdoored Models
In our main manuscript, we only presented decision boundaries for a few examples of backdoored models due to space constraints. As shown in Fig. 19d, Fig. 20d, Fig. 21, Fig. 22d, Fig. 23, Fig. 24, Fig. 25 and Fig. 26, we showcase decision boundaries for diverse backdoored models with different attack target labels (from label: 0 to label: 9 in all the datasets), including different datasets and architectures involved in our evaluations. As we can see, by arbitrarily selecting the attack target label, the corresponding target label significantly influences the decision region that the prediction distribution within the decision boundary of the backdoored model is more gathered, and triple samples are encircled by a large area of the target label.
Figure 19. More visual examples of decision boundaries for backdoored models in CIFAR-10.
(a) The color-class schema for CIFAR-10
(b) CIFAR-10 BadNets
(c) CIFAR-10 SSBA
(d) CIFAR-10 LF
Figure 20. More visual examples of decision boundaries for backdoored models in CIFAR-10.
(a) CIFAR-10 BPP
(b) CIFAR-10 TrojanNN
(c) CIFAR-10 LIRA
(d) CIFAR-10 Blind
Figure 21. More visual examples of decision boundaries for backdoored models in ImageNet-10.
(a) The color-class schema for ImageNet-10
(b) ImageNet-10 BadNets
(c) ImageNet-10 SSBA
(d) ImageNet-10 LF
Figure 22. More visual examples of decision boundaries for backdoored models in ImageNet-10.
(a) ImageNet-10 BPP
(b) ImageNet-10 TrojanNN
(c) ImageNet-10 LIRA
(d) ImageNet-10 Blind
Figure 23. More visual examples of decision boundaries for backdoored models in GTSRB.
(a) The color-class schema for GTSRB
(b) GTSRB BadNets
(c) GTSRB SSBA
(d) GTSRB LF
Continued on next page.
Figure 24. More visual examples of decision boundaries for backdoored models in GTSRB.
(a) GTSRB BPP
(b) GTSRB TrojanNN
(c) GTSRB LIRA
(d) GTSRB Blind
Continued from previous page.
Figure 25. More visual examples of decision boundaries for backdoored models in CIFAR100.
(a) The color-class schema for CIFAR100
(b) CIFAR100 BadNets
(c) CIFAR100 SSBA
(d) CIFAR100 LF
Figure 26. More visual examples of decision boundaries for backdoored models in CIFAR100.
(a) CIFAR100 BPP
(b) CIFAR100 TrojanNN
(c) CIFAR100 LIRA
(d) CIFAR100 Blind